OpenStudy (anonymous):

WIL GIVE MEDAL The sum of two numbers is 62. The larger number is 8 more than the smaller number. What are the numbers?

OpenStudy (anonymous):

@Tommynaiter

OpenStudy (mathmate):

Sum and difference problems. ================== Hints: The larger number = (sum+difference)/2 The smaller number=(sum-difference)/2

OpenStudy (anonymous):

Thank you!

OpenStudy (texaschic101):

L = larger number S = smaller number L + S = 62 L = S + 8 now you can solve by substitution